OpenStudy (anonymous):

Solving systems story problems... 6 grapefruit cost as much as a dozen oranges. The cost of a dozen grapefruit and two dozen oranges is $12. How much does 1 grapefruit and one orange cost?

OpenStudy (radar):

Let x = cost of one grape fruit. then (1/2)x= cost of one orange. Now 12x +24(1/2)x =$12.00 12x +12x =$12.00 24x=$12 x = .50 (fifty cents or .5 dollar) cost of one grapefruit x/2=50/2 = .25 (quarter) for one orange. Verify: 12 (.5) + 24 (.25) = $12 6 + 6 = 12 12=12 checks out